OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y= ((2)/(x^2))+3

OpenStudy (anonymous):

the end of f(x) would have +3...so would f(x) be (1/x^2) +3

OpenStudy (anonymous):

you have lots of ways to do this

OpenStudy (anonymous):

ok...well what i wrote.. is it right?

OpenStudy (anonymous):

it is \[y=\frac{2}{x^2}+3\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you need to write both an \(f\) and a \(g\) to answer this

OpenStudy (anonymous):

i know but does my f fit?

OpenStudy (anonymous):

here is one example \[g(x)=x^2,f(x)=\frac{2}{x}+3\]

OpenStudy (anonymous):

you will need to put a \(+3\) at the end of \(f\) yes, that is a must

OpenStudy (anonymous):

as a matter of fact, you could just write \(f(x)=x+3\) and then \(g(x)=\frac{2}{x^2}\)

OpenStudy (anonymous):

I understand! so another answer may be, f(x)=x+3 and- ...iw as about to say that

OpenStudy (anonymous):

yeah that would work lol

OpenStudy (anonymous):

haha XD thanks

OpenStudy (anonymous):

yw